# Strategic Classification with Non-Linear Classifiers

### v1

@@ 6.2 Universality and Approximation
Our results in Sec. 5.3 indicate four "types" of decision boundaries that cannot be attributed to any effective classifier, though others may exist. This implies that several basic classifiers are not realizable in the strategic setting and cannot be approximated well; a simple example is one that classifies some $B_{0.9\alpha}(x)$ as positive and is negative elsewhere. This drives our principle conclusion:
@@
Common examples of universal approximator classes include polynomial thresholds [30], RBF kernel machines [26], many classes of neural networks [e.g., 17], and gradient boosting machines [13]. In the strategic setting, these classes are no longer all-encompassing like in the standard setting.
@@
Cor. 2 implies that even if the data is in itself realizable, and even without any limitations on the hypothesis class, the learner may already start off with a bounded maximum training accuracy simply by learning in the strategic setting; see example in Fig. 5 (right) for an example dataset with limited maximum accuracy. This is because strategic behavior makes it impossible to generally approximate intricate functions, e.g., by using fast-changing curvature (via increased depth) or many piecewise-linear segments (with ReLUs). Another result is that the common practice of interpolating the data using highly expressive over-parametrized models (e.g., as in deep learning) is no longer a viable approach. On the upside, it may be possible for the strategic responses to prevent unintentional overfitting, allowing for better generalization.
In extreme cases, the gap between the maximum standard accuracy and the maximum strategic accuracy can be quite large. Our next result demonstrates that there are distributions in which despite a maximum standard accuracy of 1, the maximum strategic accuracy falls to the majority class rate.
@@
For all $d$, there exist (non-degenerate) distributions on $\mathbb{R}^{d}$ that are realizable in the standard setting, but whose strategic accuracy is the majority class rate, $\max_{y}p(y)$.
In one dimension, the distribution can be constructed by placing a negative point $\delta$ to the left and right of each positive point. In higher dimensions, negative points can be arranged in a $\mathbb{R}^{d}$ tetrahedron around each positive point. The full proof is in Appendix  B.11. Prop. 6 extends this idea to show that there is a distribution whose maximum strategic accuracy approaches $\nicefrac{{1}}{{2}}$ as $\alpha$ increases. The general proof is given in Appendix B.12.
@@
There exists a distribution which is realizable in the standard setting, but whose accuracy under any $h_{\Delta}$ is at most $0.5+\frac{1}{\lfloor 2\alpha+1\rfloor}$.
@@ 7 Experiments
To complement our theoretical results, we present two experiments that empirically demonstrate the effects of strategic behavior on non-linear classifiers.
7.1 Expressivity
Following our results from Sec. 6.1 on the change in hypothesis class expressivity, we experimentally test whether the $h_{\Delta}$ of a given random $h$ belongs to a more or less expressive class than $h$ itself. In particular, we focus on $k$-degree polynomial classifiers and compare the degree $k$ of $h$ to the degree $k_{\Delta}$ of the polynomial approximation of $h_{\Delta}$. [^3] Though the $h_{\Delta}$ of a $k$-degree polynomial $h$ is not necessarily a polynomial, we can compare the expressivity of $h$ and $h_{\Delta}$ by finding the lowest $k_{\Delta}$-degree polynomial $g$ that well-approximates $h_{\Delta}$. Because user features are often naturally bounded, we assume that $x\in{\cal{X}}=R^{2}$ and that $x$ is bounded by $|x|_{\infty}\leq 100$. See Appendix D.1 for full details.
Fig. 6 (left) shows results for varying $k\in[1,10]$ and for $\alpha\in[2,40]$. When $\alpha$ is small, we find that $k_{\Delta}\approx k$, meaning that strategic behavior has little impact on expressivity. However, as $\alpha$ increases, $k_{\Delta}$ becomes larger that $k$, suggesting that $h_{\Delta}$ is more complex than $h$. This is due to the fact that as $\alpha$ increases, point mapping collisions (see Sec. 4) become more likely causing non-smooth cusps – which are more complex than basic polynomials – to form in the decision boundary (Fig. 6 (center)). When $\alpha$ is quite large, we see that $k_{\Delta}$ is still larger for low $k$, but drops considerably for higher $k$. This can be attributed to the fact that tightly-embedded higher-dimension polynomials and increased strategic reach from large $\alpha s$ are both causes of indirect wipeout (see Sec. 4) because they increase the reachability of $\nabla_{h}(x)$ by other decision boundary points. As such, much of the original decision boundary is wiped out leaving behind a lower complexity effective decision boundary.
7.2 Approximation
As seen in Sec. 6.2, the lack of universality in the strategic setting can impose a limitation on the maximal attainable strategic accuracy. To demonstrate the practical implications of this effect, we upper bound the maximum strategic accuracy of any $H_{\Delta}$ on a set of synthetic data by experimentally calculating the maximum strategic accuracy of the unrestricted effective hypothesis class $H^{*}_{\Delta}$. We compare the results to two benchmarks: (i) the standard accuracy of the regular unrestricted hypothesis class $H^{*}$ (which is always 1), and (ii) the strategic (and standard) accuracy of the linear effective hypothesis class $H^{\mathrm{lin}}_{\Delta}=H^{\mathrm{lin}}$. The first benchmark measures the amount to which the strategic setting hinders the learner, while the second measures the amount to which the learner can benefit from non-linearity, even in the strategic world. We generate data by sampling points from class-conditional Gaussians $x\sim{\cal{N}}(y\mu,1)$, and aim to learn an $h$ that obtains an optimal fit. When $\mu$ is large, the classes are likely to be well-separated, but when $\mu$ is small, points from different classes will be more "interleaved" and thus harder to classify by restricted classes like $H^{\mathrm{lin}}_{\Delta}$ and $H^{*}_{\Delta}$. Details in Appendix D.1.
Fig. 6 (right) shows results for increasing class separation ($\mu$). As $\mu$ decreases and the classification problem becomes more difficult, the maximum strategic accuracy of $H^{*}_{\Delta}$ diverges from the upper $H^{*}$ baseline, with performance deteriorating as strategic behavior intensifies (i.e., larger $\alpha$). Because any actual $H_{\Delta}$ can only have worse performance than the ideal $H^{*}_{\Delta}$, this indicates that strategic behavior can become a significant burden in more difficult tasks such as classification in a non linearly-separable setting. Nonetheless, when $\mu$ is small, $H^{*}_{\Delta}$ significantly outperforms the lower $H^{\mathrm{lin}}_{\Delta}$ baseline signaling that non-nonlinearities still improve accuracy in the strategic setting despite any strategic effects.
@@ 8 Discussion
This work sets out to explore the interplay between non-linear classifiers and strategic user behavior. Our analysis demonstrates that non-linearity induces behavior that is both qualitatively different than the linear case and also non-apparent by simply studying it. Simply put, the strategic setting is fundamentally limited, though has a few potential advantages. Our results show how such behavior can impact classifiers by morphing the decision boundary and model classes by increasing or decreasing complexity. Although prior work has made clear the importance of accounting for strategic behavior in the learning objective, our work suggests that there is a need for additional broader considerations throughout the entire learning pipeline from the choices we make initially (such as which model class to use) to setting our final expectations (such as the accuracy levels we may aspire to achieve). This motivates future theoretical questions, as well as complementary work on practical aspects such as optimization (i.e., how to solve the learning problem), modeling (i.e., how to capture true human responses), and evaluation (i.e., on real humans and in the wild).
Appendix A Definitions
@@ A.1 Signed Curvature
Our analysis makes use of the notion of signed curvature – a signed version of the normal curvature – to quantify shape alteration. Because there is an inherent notion of direction due to the binary nature of label classes, we can define curvature as being positive if the curvature normal vector points in the direction of the positive region and negative if it points in the direction of the negative region. Though in $\mathbb{R}^{2}$ this defines a single value, we must instead define a curvature value for each "direction" in higher dimensions. More formally, for each tangent vector $\vec{t}$ in the tangent hyperplane $T_{x}$ at $x$, we define the curvature value $\kappa_{\vec{t}}$ as the signed curvature at $x$ on the curve obtained by intersecting the decision boundary with the plane $L$ containing $\vec{t}$ and the normal vector $\hat{n}_{x}$. Defining $K=\{\kappa_{\vec{t}}:\vec{t}\in T\}$ as the set of all such curvatures, our analysis will focus on $\kappa^{*}=\sup(K)$ in $\mathbb{R}^{3}$ and above.
@@ Appendix B Proofs
In $\mathbb{R}^{2}$, if the curvature $\kappa$ at $x$ is positive, then the radius $r_{insc}$ of the maximum inscribed circle at $x$ is no bigger than the radius $r_{osc}=\frac{1}{\kappa}$ of the osculating circle at $x$ (the circle that best approximates the curve at $x$ and has the same curvature). As a corollary, if curvature $\kappa$ at $x$ is negative, then the radius $r_{esc}$ of the maximum escribed circle at $x$ is no bigger than the radius $r_{osc}=\frac{1}{-\kappa}$ of the osculating circle at $x$.
@@
Note that both the osculating circle and the inscribed circle have their centers along the normal line to the curve at $x$ (and on the same side of $h$), and that the osculating circle may cross the curve at $x$. If it does, then any circle which has a larger radius and a center on the same side of $h$ along the normal line to the curve will also cross the decision boundary at $x$, so the radius of the inscribed circle is strictly smaller than the radius of the osculating circle. If the osculating circle does not cross the curve, then it must be the largest strictly tangent circle at $x$ or else it would not be the best circular approximation to the curve at $x$. As such, the largest inscribed circle, which must be both tangent to the curve at $x$ also not intersect it anywhere else, must have a radius that is no larger than $r_{osc}$.
@@ B.1 One-to-One Mapping
In one-to-one mapping, the point $x$ on the decision boundary of $h$ is mapped to the point ${\hat{x}}=x-\alpha\hat{n}_{x}$ on the decision boundary of $h_{\Delta}$.
@@
Note that all points on the decision boundary of $h_{\Delta}$ must be at the maximum cost away from $h$ [^4] Otherwise, $x$ would not be on the on the boundary of positive and negative classification by $h_{\Delta}$ because all of the points adjacent to it can also strategically move to get a positive classification., so $\nabla_{h}(x)\subset S_{\alpha}(x)$. Additionally, for each $x^{\prime}\in\nabla_{h}(x)$, the ball $B_{\alpha}(x^{\prime})$, which is the set of points that $x^{\prime}$ can move to, must not intersect the decision boundary of $h$ (or else it would not be at maximum cost away) and must furthermore be strictly tangent to the decision boundary at $x$. Given that the decision boundary of $h$ is smooth at $x$, this means that the normal to the curve at $x$ must run along the radius of $B_{\alpha}(x^{\prime})$ from $x$ to $x^{\prime}$, implying that $x^{\prime}=x\pm\alpha\hat{n}_{x}$. Because of the inherit directionality of the setting, the normal vector at $x$ is defined to point in the positive direction. This means that only the point ${\hat{x}}=x-\alpha\hat{n}_{x}$ would seek to move to $x$ to gain a positive classification as $x^{\prime}=x+\alpha\hat{n}_{x}$ is either already labeled positive or can move to a closer point. Therefore, $\nabla_{h}(x)=\{x-\alpha\hat{n}_{x}\}$ in the one-to-one case. ∎
@@ B.2 Wipeout Mapping
Points on the decision boundary of $h$ that are "wiped out" do not affect the effective decision boundary of $h_{\Delta}$.
@@
In the case of indirect wipeout, this follows directly from the definition: if $\nabla_{h}(x)$ is contained in $B_{\alpha}(x^{\prime})$ (or in the union of several $B_{\alpha}(x^{\prime})$), then none of the points in $\nabla_{h}(x)$ are at a maximum cost away from $h$ and therefore are not on the decision boundary of $h_{\Delta}$.
In the case of direct wipeout, the normal one-to-one mapping fails because the point ${\hat{x}}=x-\alpha\hat{n}_{x}$ is no longer at the maximum cost away from $h$. Specifically, because the curvature $\kappa<-1/\alpha$ 555or $\inf(K)<-1/\alpha$ for $d>2$. See Sec. 5.1 and Appendix A.1 for curvature definitions, the ball $B_{\alpha}(x^{\prime})$ now intersects $h$, which means that $x^{\prime}$ is less than $\alpha$ away from $h$. For $\mathbb{R}^{2}$, Lemma 1 proves that the radius of the maximum escribed circle at $x$ is less than the radius of the osculating circle, so $r_{esc}\leq r_{osc}=\frac{1}{-\kappa}<\alpha$. Therefore, $B_{\alpha}(x^{\prime})$ must intersect $h$.
For higher dimensions (see Appendix A.1), there is no longer a single normal curvature. Instead, we define the set $K$ of curvatures in each plane $p\in P$ defined by the normal vector and some tangent vector at $x$. Note that $B_{\alpha}(x^{\prime})$ will intersect $h$ if any of its projections onto each $p\in P$ (a circle) intersects the curve obtained by intersecting the decision boundary with the plane $p$. Since each of these cases are in $\mathbb{R}^{2}$, $B_{\alpha}(x^{\prime})$ will intersect $h$ if $\inf(K)<-1/\alpha$.

### v2

@@ 5.2 Universality and approximation
Our results in Sec. 4.4 show four ‘types’ of decision boundaries that cannot be attributed to any effective classifier, though others may exist. This implies that several basic classifiers are not realizable in the strategic setting and cannot be approximated well. This drives our principal conclusion:
@@
In other words, there exists a classifier $g$ for which no effective class $H_{\Delta}$ can include $g$ as a member, even for $H$ that are universal approximators in the non-strategic setting. Common examples of universal approximator classes include polynomial thresholds [32], RBF kernel machines [27], many classes of neural networks [e.g., 17], and gradient boosting machines [13]. In the strategic setting, these classes are no longer all-encompassing like in the standard setting.
@@
Cor. 2 implies that even if the data is in itself realizable, and even without any limitations on the hypothesis class, the learner may already start off with a bounded maximum training accuracy simply by learning in the strategic setting. Fig. 4 (right) depicts one such example dataset where the maximum attainable accuracy drops from 1 to 0.96. This is because strategic behavior makes it impossible to generally approximate intricate functions, e.g., by using fast-changing curvature (via increased depth) or many piecewise-linear segments (with ReLUs). Another result is that the common practice of interpolating the data using highly expressive over-parametrized models (e.g., as in deep learning) is no longer a viable approach. On the upside, it may be possible for the strategic responses to prevent unintentional overfitting, allowing for better generalization.
In extreme cases, the gap between the maximum standard accuracy and the maximum strategic accuracy can be quite large. Our next result demonstrates that there are distributions in which, despite a maximum standard accuracy of 1, the maximum strategic accuracy falls to the majority class rate.
@@
For all $d$, there exist (non-degenerate) distributions on $\mathbb{R}^{d}$ that are realizable in the standard setting, but whose maximum strategic accuracy is the majority class rate, $\max_{y}p(y)$.
In one dimension, the distribution can be constructed by placing a negative point $\delta$ to the left and right of each positive point. In higher dimensions, negative points can be arranged in an $\mathbb{R}^{d}$ simplex around each positive point. The full proof is in Appendix  B.12. Prop. 6 extends this idea to show that there is a distribution whose maximum strategic accuracy approaches $\nicefrac{{1}}{{2}}$ as $\alpha$ increases. The general proof is given in Appendix B.13.
@@
There exists a distribution that is realizable in the standard setting, but whose accuracy under any $h_{\Delta}$ is at most $0.5+\frac{1}{\lfloor 2\alpha+1\rfloor}$.
@@ 6 Experiments
To complement our theoretical results, we present two experiments that empirically demonstrate the effects of strategic behavior on non-linear classifiers. The code for both experiments is available at https://github.com/BML-Technion/scnonlin.
@@
6.1 Expressivity
Following our results from Sec. 5.1 on the change in hypothesis class expressivity, we experimentally test whether the $h_{\Delta}$ of a given random $h$ belongs to a more or less expressive class than $h$ itself. In particular, we focus on degree-$k$ polynomial classifiers and compare the degree $k$ of $h$ to the degree $k_{\Delta}$ of the polynomial approximation of $h_{\Delta}$. [^2] Though the $h_{\Delta}$ of a polynomial $h$ is not necessarily a polynomial, we can compare the expressivity of $h$ and $h_{\Delta}$ by finding the lowest degree-$k_{\Delta}$ polynomial $g$ that well-approximates $h_{\Delta}$. Because user features are often naturally bounded, we assume that $x\in{\cal{X}}=R^{2}$ and that $x$ is bounded by $|x|_{\infty}\leq 100$. See Appendix D.1 for full details.
Fig. 5 (left) shows results for varying $k\in[1,10]$ and for $\alpha\in[2,40]$. When $\alpha$ is small, we find that $k_{\Delta}\approx k$, meaning that strategic behavior has little impact on expressivity. However, as $\alpha$ increases, $k_{\Delta}$ becomes larger than $k$, suggesting that $h_{\Delta}$ is more complex than $h$. This is due to the fact that as $\alpha$ increases, point mapping collisions (see Sec. 4.1) become more likely, causing non-smooth cusps — which are more complex than basic polynomials — to form in the decision boundary (Fig. 5 (center)). When $\alpha$ is quite large, we see that $k_{\Delta}$ is still larger for low $k$, but drops considerably for higher $k$. This can be attributed to the fact that tightly-embedded higher-dimension polynomials and increased strategic reach from large $\alpha s$ are both causes of indirect wipeout (see Sec. 4.1) because they increase the reachability of $\nabla_{h}({z})$ by other decision boundary points. As such, much of the original decision boundary is wiped out, leaving behind a lower complexity effective decision boundary.
6.2 Approximation
As seen in Sec. 5.2, the lack of universality in the strategic setting can impose a limitation on the maximal attainable strategic accuracy. To demonstrate the practical implications of this effect, we upper bound the maximum strategic accuracy of any $H_{\Delta}$ on a set of synthetic data by experimentally calculating the maximum strategic accuracy of the unrestricted effective hypothesis class $H^{*}_{\Delta}$. We compare the results to two benchmarks: (i) the standard accuracy of the regular unrestricted hypothesis class $H^{*}$ (which is always 1), and (ii) the strategic (and standard) accuracy of the linear effective hypothesis class $H^{\mathrm{lin}}_{\Delta}=H^{\mathrm{lin}}$. The first benchmark measures the extent to which the strategic setting hinders the learner, while the second measures the extent to which the learner can benefit from non-linearity, even in the strategic world. We generate data by sampling points from class-conditional Gaussians $x\sim{\cal{N}}(y\mu,1)$, and aim to find an $h$ that obtains an optimal fit. The parameter $\mu$ serves to show how $H_{\Delta}$ compares to our two baselines under data that ranges from well-separated (large $\mu$) to more "interleaved" and harder to classify by restricted classes (small $\mu$). Details in Appendix D.2.
Fig. 5 (right) shows results for increasing class separation ($\mu$). As $\mu$ decreases and the classification problem becomes more difficult, the maximum strategic accuracy of $H^{*}_{\Delta}$ diverges from the upper $H^{*}$ baseline, with performance deteriorating as strategic behavior intensifies (i.e., larger $\alpha$). Because any actual $H_{\Delta}$ can only have worse performance than the ideal $H^{*}_{\Delta}$, this indicates that strategic behavior can become a significant burden in more difficult tasks, such as classification in a non linearly-separable setting. Nonetheless, when $\mu$ is small, $H^{*}_{\Delta}$ significantly outperforms the lower $H^{\mathrm{lin}}_{\Delta}$ baseline, signaling that non-linearities still improve accuracy in the strategic setting despite any strategic effects.
@@ 7 Discussion
This work sets out to explore the interplay between non-linear classifiers and strategic user behavior. Our analysis demonstrates that non-linearity induces behavior that is both qualitatively different than the linear case and also non-apparent by simply studying it. Simply put, the strategic setting is fundamentally limited, though has a few potential advantages. Our results show how such behavior can impact classifiers by morphing the decision boundary and model classes by increasing or decreasing complexity. Although prior work has made clear the importance of accounting for strategic behavior in the learning objective, our work suggests that there is a need for additional broader considerations throughout the entire learning pipeline from the choices we make initially (such as which model class to use) to setting our final expectations (such as the accuracy levels we may aspire to achieve). We detail practical takeaways for learning and model class selection in Appendix C.10. This motivates future theoretical questions, as well as complementary work on practical aspects such as optimization (i.e., how to solve the learning problem), modeling (i.e., how to capture true human responses), and evaluation (i.e., on real humans and in the wild).
Acknowledgments
The authors are grateful to Shay Moran and Nadav Dym for thoughtful discussions and suggestions. This work is supported by the Israel Science Foundation grant no. 278/22.
Appendix A Background
@@ A.1 Signed Curvature
Our analysis makes use of the notion of signed curvature — a signed version of the normal curvature — to quantify shape alteration. Because there is an inherent notion of direction due to the binary nature of class labels, we can define curvature as being positive if the curvature normal vector points in the direction of the positive region and negative if it points in the direction of the negative region. Though in $\mathbb{R}^{2}$ this defines a single value, we must instead define a curvature value for each ‘direction’ in higher dimensions. More formally, for each tangent vector $\vec{t}$ in the tangent hyperplane $T_{x}$ at $x$, we define the curvature value $\kappa_{\vec{t}}$ as the signed curvature at $x$ on the curve obtained by intersecting the decision boundary with the plane $L$ containing $\vec{t}$ and the normal vector $\hat{n}_{x}$. Defining $K=\{\kappa_{\vec{t}}:\vec{t}\in T\}$ as the set of all such curvatures, our analysis will focus on $\kappa^{*}=\sup(K)$ in $\mathbb{R}^{3}$ and above.
A.2 Bounded/Unbounded Input Domain
At a high level, our results in Sec. 5 do not require an unbounded input space, but rather some input space that is larger than $\cal{X}$}. In particular, if we assume that raw inputs, $x$, are bounded by $|x|\leq B$, then for the problem to generally be well-defined, we must allow this space to expand at the minimum to $B+\alpha$ to accommodate for strategic input modifications. More generally, we must ensure that $|\Delta_{h}(x)|\leq B(\alpha)$ holds for all $x$ and any $h$, for some expansion $B(\alpha)\geq B$, where the exact form of $B(\alpha)$ depends both on the cost function and on $H$.
Other than Thm. 2, our results simply require the minimal expansion be satisfied and are therefore not dependent on an unbounded input domain. Likewise, Thm. 2 does not require an unbounded input domain since for any given learnable $H$, there exists some finite $B(\alpha)$ for which the result holds. However, since the statement is generic (i.e., applies to all learnable $H$ that are closed to scaling), we do not have a closed-form expression for $B(\alpha)$ that applies simultaneously to all such $H$.
A.3 General Upper Bound on VC Increase
In Thm. 3, we provide an upper bound on the strategic VC dimension of classes of piecewise-linear classifiers, though note that a general upper bound remains an open question. A key difficulty in finding a general upper bound in the strategic setting is that points will move differently when facing different $h$ from the same $H$. Thus, reasoning about the shattering of an arbitrary set must account for how points can move under all possible $h\in H$, though not simultaneously, which proved to be highly challenging in the general, unstructured case. This is why we have included results that exploit common structures like the upper bound based on piecewise linearity (Thm. 3) and lower bound based on $H$ that are closed to scaling (Thm. 2).
That said, we conjecture that Thm. 3 holds for general $H$, as proven by Cohen et al. [6] in the special case of finite manipulation graphs. In particular, they show that $\mathrm{VC}(H_{\Delta})=\tilde{\Theta}(\mathrm{VC}(H)\cdot log\,k)$ where $k$ is the degree of the manipulation graph. Note that the bound is not transferable to our case since it is vacuous for $k\to\infty$, and the proof technique does not apply to continuous graphs.
@@ Appendix B Proofs
In $\mathbb{R}^{2}$, if the curvature $\kappa$ at $x$ is positive, then the radius $r_{insc}$ of the maximum inscribed circle at $x$ is no bigger than the radius $r_{osc}=\frac{1}{\kappa}$ of the osculating circle at $x$ (the circle that best approximates the curve at $x$ and has the same curvature). As a corollary, if curvature $\kappa$ at $x$ is negative, then the radius $r_{esc}$ of the maximum escribed circle at $x$ is no bigger than the radius $r_{osc}=-\frac{1}{\kappa}$ of the osculating circle at $x$.
@@
Note that both the osculating circle and the inscribed circle have their centers along the normal line to the curve at $x$ (and on the same side of $h$), and that the osculating circle may cross the curve at $x$. If it does, then any circle which has a larger radius and a center on the same side of $h$ along the normal line to the curve will also cross the decision boundary at $x$, so the radius of the inscribed circle (which strictly does not cross the curve) is strictly smaller than the radius of the osculating circle. If the osculating circle does not cross the curve, then it must be the largest strictly tangent circle at $x$, or else it would not be the best circular approximation to the curve at $x$. As such, the largest inscribed circle, which must be both tangent to the curve at $x$ also not intersect it anywhere else, must have a radius that is no larger than $r_{osc}$.
@@ B.1 One-to-One Mapping
In the one-to-one mapping, the point z{z} on the decision boundary of $h$ is mapped to the point ${x}={z}-\alpha\hat{n}_{z}$ on the decision boundary of $h_{\Delta}$.
@@
Note that all points on the decision boundary of $h_{\Delta}$ must be at the maximum cost away from hh [^3] Otherwise, $x$ would not be on the boundary of positive and negative classification by $h_{\Delta}$ because all of the points adjacent to it can also strategically move to achieve a positive classification., so $\nabla_{h}({z})\subset S_{\alpha}({z})$. Additionally, for each ${x}\in\nabla_{h}({z})$, the ball $B_{\alpha}({x})$, which is the set of points that x{x} can move to, must not intersect the decision boundary of $h$ (or else it would not be at maximum cost away) and must furthermore be strictly tangent to the decision boundary of $h$ at z{z}. Given that the decision boundary of $h$ is smooth at z{z}, this means that the normal to the curve at z{z} must run along the radius of $B_{\alpha}({x})$ from z{z} to x{x}, implying that ${x}={z}\pm\alpha\hat{n}_{z}$. Because of the inherent directionality of the setting, the normal vector at z{z} is defined to point in the positive direction. This means that only the point ${x}={z}-\alpha\hat{n}_{z}$ would seek to move to z{z} to gain a positive classification, as ${x}={z}+\alpha\hat{n}_{z}$ is either already labeled positive or can move to a closer point. Therefore, $\nabla_{h}({z})=\{{z}-\alpha\hat{n}_{z}\}$ in the one-to-one case. ∎
@@ B.2 Wipeout Mapping
Points on the decision boundary of $h$ that are ‘wiped out’ do not affect the effective decision boundary of $h_{\Delta}$.
@@
In the case of indirect wipeout, this follows directly from the definition: if $\nabla_{h}({z})$ is contained in $B_{\alpha}(x^{\prime})$ for some other $x^{\prime}$ on the decision boundary of $h$ (or in the union of several $B_{\alpha}(x^{\prime})$), then none of the points in $\nabla_{h}({z})$ are at a maximum cost away from $h$, and are therefore not on the decision boundary of $h_{\Delta}$.
In the case of direct wipeout, the normal one-to-one mapping fails because the point ${x}={z}-\alpha\hat{n}_{z}$ is no longer at the maximum cost away from $h$. Specifically, because the curvature $\kappa<-1/\alpha$ 444or $\inf(K)<-1/\alpha$ for $d>2$. See Sec. 4.2 and Appendix A.1 for curvature definitions, the ball $B_{\alpha}({x})$ now intersects $h$, which means that x{x} is less than $\alpha$ away from $h$ and therefore not on the decision boundary of $h_{\Delta}$. For $\mathbb{R}^{2}$, Lemma 1 proves that the radius of the maximum escribed circle at z{z} is less than the radius of the osculating circle, so $r_{esc}\leq r_{osc}=-\frac{1}{\kappa}<\alpha$. Therefore, there is no circle of radius $\alpha$ that is strictly tangent to $h$ (on the negative side), so $B_{\alpha}({x})$ must intersect $h$.
For higher dimensions (see Appendix A.1), there is no longer a single normal curvature. Instead, we define the set $K$ of curvatures in each plane $p\in P$ defined by the normal vector and some tangent vector at z{z}. Note that $B_{\alpha}({x})$ will intersect $h$ if any of its projections onto each $p\in P$ (a circle) intersects the curve obtained by intersecting the decision boundary with the plane $p$. Since each of these cases are in $\mathbb{R}^{2}$, $B_{\alpha}({x})$ will intersect $h$ if $\inf(K)<-1/\alpha$.